An official of a large national union claims that the fraction of women in the union is not significantly different from 58%. To test this claim, a random sample of 400 individuals in the union is taken. Of those 400, 215 are women.

Use the p-value to conduct the hypothesis test and use α = 0.05 level of significance.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p = 0.58

Ha : p \neq 0.58

n = 400

x = 215

\hat p = x / n = 215 / 400 = 0.54

P0 = 0.58

1 - P0 = 1 - 0.58 = 0.42

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

=0.54 - 0.58 / [\sqrt(0.58*0.42) /400 ]

= -1.72

Test statistic = z = -1.72

P(z < -1.72 ) = 0.0427

P-value = 2 * 0.0427 = 0.0854

\alpha = 0.05

P-value > \alpha

0.0854 > 0.05

Fail to reject the null hypothesis .

There is not sufficient evidence to suggest that